# Relative Entropy Methods for Calculating Committors

### v1

@@ 4.2. Selection: Estimating Relative Entropy Differences
$\tilde{\mu}:=\int_{\partial A}\tilde{m}(x)\,\mathrm{d}S_{A}(x)\text{ and }\bar{\mu}:=\int_{\partial A}\bar{m}(x)\,\mathrm{d}S_{A}(x).$
Assume that $\tilde{m}$ and $\bar{m}$ are strictly positive. We define $\tilde{\mathds{P}}$ to be the law of the approximate transition path process corresponding to $\tilde{q}$ and $\tilde{m}$, i.e. the law of a weak solution of (7) with singular drift $2\varepsilon\nabla\log\tilde{q}(Y_{t})$ and initial condition $Y_{0}\sim\tilde{\mu}^{-1}m(x)\,\mathrm{d}S_{A}(x)$. We define $\bar{\mathds{P}}$ similarly, but with $\bar{q}$ and $\bar{m}$ in place of $\tilde{q}$ and $\tilde{m}$. We now explain how to estimate the relative entropy difference
@@
Suppose that we have an independent sample of $\tilde{N}$ paths $\tilde{Y}^{k}_{t}$ drawn from $\tilde{\mathds{P}}$:
@@
$\tilde{I}$ $:=\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\mu}\,\mathrm{d}S_{A}(x)$
$\qquad+\mathds{P}_{\theta}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$
appearing in formula (12) for $D_{\mathrm{KL}}(\tilde{\mathds{P}}\|\mathds{Q})$. We give conditions on $\tilde{q}$ in Lemma 8 in Appendix B that guarantee finite variance and a weak law of large numbers. Of course, one could estimate the analogous term $\bar{I}$ in $D_{\mathrm{KL}}(\bar{\mathds{P}}\|\mathds{Q})$ by a similar average $\bar{I}_{\bar{N}}$ given a sample of $\bar{N}$ paths $\bar{Y}^{k}_{t}$ drawn from $\bar{\mathds{P}}$. The crucial problem is to show that the variance of $\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(\tilde{Y}^{k}_{s})\,\mathrm{d}s$ is finite even though $\frac{L\tilde{q}}{\tilde{q}}$ would typically be singular on $\partial A$, since $\tilde{q}=0$ on $\partial A$.
@@
$\delta(\tilde{q},\tilde{m};\bar{q},\bar{m})=\log\left(\frac{\tilde{\mu}}{\bar{\mu}}\right)+\tilde{I}-\bar{I},$
so it remains to estimate the ratio $\frac{\tilde{\mu}}{\bar{\mu}}$. Analogous problems involving ratios of normalizing constants arise in Bayesian model selection and the calculation of alchemical free energy differences. Refer to [11] for a survey of methods for free energy calculations. We propose to use the Bennett Acceptance Ratio (BAR) method [1] to estimate $\frac{\tilde{\mu}}{\bar{\mu}}$. BAR is in some sense the optimal estimator of $\frac{\tilde{\mu}}{\bar{\mu}}$ given i.i.d. samples from $\tilde{m}\,\mathrm{d}S$ and $\bar{m}\,\mathrm{d}S$. More complex alternatives that involve sampling from multiple distributions, such as the Multistate Bennett Acceptance Ratio (MBAR) method, may be needed in cases where $\tilde{m}$ and $\bar{m}$ differ greatly [15].
Remark 5.
In practice, to sample from the approximate reactive flux densities $\tilde{m}$ and $\bar{m}$, one would most likely use a Markov chain Monte Carlo (MCMC) method. Many methods have been developed to sample distributions supported on submanifolds such as $\partial A$. For example, see [11, Sections 3.2.3-4]. Of course, MCMC would produce correlated samples from the approximate reactive flux distributions, not independent samples as assumed above.
4.3. Training: Minimizing the Relative Entropy
@@
$m_{\theta}(x):=\mu_{\theta}^{-1}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$ (14)
@@
$\mu_{\theta}:=\int_{\partial A}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x).$
We define $\mathds{P}_{\theta}$ to be the law of a weak solution of the stochastic differential equation (7) observed up to time $\tau$ for the approximate committor function $q_{\theta}$ and initial condition $Y_{0}\sim m_{\theta}$. In this section, we derive a consistent estimator of $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ given a sample from $\mathds{P}_{\theta}$. We then use the estimator in a stochastic gradient descent method to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$.
For our results in this section, we require much stronger assumptions on $q_{\theta}$ and on the distribution of $\tau$ than those imposed elsewhere; cf. Assumption 7. The most significant of these new assumptions is that
$Lq_{\theta}=0\text{ on }\partial A\text{ for all }\theta\in\mathbb{R}^{k}.$ (15)
Recall that for the exact committor $q$, or more precisely for the smooth extension guaranteed by Lemma 1, we have $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$; cf. Remark 2. One can construct practical families $q_{\theta}$ satisfying (15) (cf. Section 5), and our numerical experiments indicate that stochastic gradient descent may converge faster when (15) holds. However, we do not claim that (15) is necessary for differentiability.
@@
that appears in the change of measure formula (11) extends to a smooth function of $x$ defined on an open neighborhood of $\partial A$, even though we assume $q_{\theta}=0$ on $\partial A$. To see this, let $Lq_{\theta}$ take the place of $\tilde{q}$ in the proof of Lemma 2. Similarly,
@@
$\mathds{P}_{\theta}[\exp(\gamma(\theta)\tau)]<\infty;$ (16)
that is, the moment generating function of $\tau$ under $\mathds{P}_{\theta}$ is finite on an open interval containing zero. We will not verify (16) for any family of approximate committor functions. However, we note that for the exact transition path process in one-dimension, under some rather stringent conditions on the potential energy $U$ and on the sets $A$ and $B$, one can show that $\tau$ converges in distribution as $\varepsilon\rightarrow 0$ to a shifted and scaled Gumbel random variable [2]. Under much more general conditions, escape times from basins of attraction of minima of $U$ are approximately exponentially distributed for the overdamped Langevin dynamics in the limit of small $\varepsilon$ [8]. The moment generating function of a Gumbel or exponential random variable will be finite on an open interval containing zero, but not globally finite, which motivates (16).
@@
    $M\geq\mu_{\theta}\geq\frac{1}{M}>0\text{ and }\lvert\nabla_{\theta}m_{\theta}\rvert\leq M$
@@
$\mathds{P}_{\theta}[\exp(\gamma(\theta)\tau)]<\infty.$ (17)
Under Assumption 7, the relative entropy is differentiable.
Theorem 2.
Let Assumption 7 hold, and let $\mathds{P}_{\theta}$ and $m_{\theta}$ be defined as above. The relative entropy $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ is differentiable as a function of $\theta$, and
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ $=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\nabla_{\theta}\log q_{\theta}(Y_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y_{s})\,\mathrm{d}s,\right.$
@@
$\begin{split}G_{N}&:=\frac{1}{N-1}\sum_{k=1}^{N}\left(\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-J_{N}\right)\\ &\qquad\qquad\qquad\times\left(\nabla_{\theta}\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-K_{N}\right).\end{split}$ (18)
Assumption 7 implies that $G_{N}$ has finite variance. We leave the proof to the reader. Using this gradient estimator, one can attempt to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ by stochastic gradient descent.
4.4. Importance Sampling and Direct Estimation of Relative Entropy
Using our change of measure formula, one could in principle estimate arbitrary expectations over the transition path process $\mathds{Q}$ given a sample of paths from the approximate process $\mathds{P}$. Let $g:\Omega\rightarrow\mathbb{R}$ be $\mathcal{F}_{\tau}$-measurable, i.e. a function of the process observed up to time $\tau$. By (11), assuming for simplicity that $\tilde{q}=1$ on $\partial B$ and $m=\lvert\nabla\tilde{q}\rvert\exp(-U/\varepsilon)$, we have
$\mathds{Q}[g]$ $=\mathds{P}\left[Z_{\tau}^{-1}g\right]$
$=\frac{\mu}{\nu}\mathds{P}\left[\exp\left(\int_{0}^{\tau(\omega)}\frac{L\tilde{q}}{\tilde{q}}(Y_{s}(\omega))\,\mathrm{d}s\right)g(\omega)\right],$
@@ 4.4. Importance Sampling and Direct Estimation of Relative Entropy
$\bar{g}=\frac{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)g(Y^{k})}{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)},$
where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. Experience with other estimators related to Girsanov’s theorem strongly suggests that such a naïve estimator will not be robust. Some successes have been attained in the importance sampling of paths of diffusion processes by roughly similar estimators [18, 4, 16]. We hope that our results will facilitate the development of analogous methods for the transition path process. However, we note that [18, 4, 16] treat only paths observed up to a fixed finite time, not up to an unbounded stopping time as in our work.
@@
$\frac{\nu}{\mu}$ $=\mathds{P}\left[\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)\right],$
@@
$\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right),$
where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. We do not expect such an estimator to be much more robust than the general importance sampling estimator.
5. Numerical Methods
@@
5.1. Representing the Committor Function
@@
appearing in Assumption 7. In our approach, one must first choose a computable function $T:\Omega\rightarrow\mathbb{R}$ with the following properties:
  • $T$ extends to a smooth function on an open neighborhood of $\overline{\Omega}$.
@@
  • $T(x)>0$ for all $x\in\Omega\setminus\partial A$.
@@
where $T$ is as above and $w:\Omega\rightarrow\mathbb{R}$ belongs to some convenient family of smooth functions. Any $\tilde{q}$ of this form satisfies Assumption 4. In particular, the normal derivative of $\tilde{q}$ is positive over $\partial A$, since
$\nabla\tilde{q}(x)\cdot n(x)$ =$=\big{\{}\nabla T(x)\exp(w(x))+\nabla w(x)T(x)\exp(w(x))\big{\}}\cdot n(x)$
@@
for $x\in\partial A$. Moreover, for any $T$ meeting the conditions above, the exact committor $q$ can be written in the form $q=T\exp(w)$ for $w=\log\left(q/T\right)$, and this $w$ extends to a smooth function on an open neighborhood of $\Omega$ by Lemma 2.
The condition $L\tilde{q}=0$ on $\partial A$ appearing in Assumption 7 is equivalent with a Neumann boundary condition on $w$. We have
$L\tilde{q}=\left\{LT+2\varepsilon\nabla w\cdot\nabla T+T\left(Lw+\varepsilon\lvert\nabla w\rvert^{2}\right)\right\}\exp(w).$ (19)
@@
$\nabla w(x)\cdot n(x)=-\frac{1}{2\varepsilon}\frac{LT(x)}{\lvert\nabla T(x)\rvert}$ (20)
for $x\in\partial A$.
Depending on the geometry of $\partial A$, it may be more or less difficult to impose the Neumann boundary condition. We propose a general approach that requires a smooth global retraction $\rho:\Omega\rightarrow\partial A$ in addition to $T$. We represent $w$ in the form
$w(x)=w_{0}(\rho(x))+T(x)w_{1}(\rho(x))+T(x)^{2}w_{2}(x).$ (21)
Here, $w_{0}:\partial A\rightarrow\mathbb{R}$ and $w_{2}:\Omega\rightarrow\mathbb{R}$ are arbitrary. Given $w_{0}$, the condition $L\tilde{q}=0$ on $\partial A$ determines $w_{1}:\partial A\rightarrow\mathbb{R}$. See Lemma 6 below for details.
Lemma 6.
Any function $w:\Omega\rightarrow\mathbb{R}$ that admits a smooth extension to an open neighborhood of $\overline{\Omega}$ may be expressed in the form (21) for some unique smooth $w_{0}$, $w_{1}$, and $w_{2}$. Moreover, the approximate committor function $\tilde{q}=T\exp(w)$ has $L\tilde{q}=0$ on $\partial A$ if and only if
$w_{1}(z)=-\frac{\nabla w_{0}^{t}(z)D\rho(z)n(z)}{\lvert\nabla T(z)\rvert}-\frac{LT(z)}{\lvert\nabla T(z)\rvert^{2}}$ (22)
for all $z\in\partial A$. In particular, the exact committor function $q$ can be expressed as $q=T\exp(w)$ with $w$ of the form (21) for some smooth $w_{0}$, $w_{1}$, and $w_{2}$.
Proof.
See Appendix C. ∎
Observe that $w_{0}$ determines the approximate reactive flux distribution corresponding to $\tilde{q}$. When $w$ takes the form (21), we have
$\nabla\tilde{q}(x)=\nabla T(x)\exp(w_{0}(x))$
for $x\in\partial A$, and therefore
$\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x)=\lvert\nabla T(x)\rvert\exp(w_{0}(x))\exp(U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
using that $\nabla T=\lvert\nabla T\rvert n$ on $\partial A$, as explained above.
5.2. Alternative Expression for the Improper Integral
Observe that $\frac{L\tilde{q}}{\tilde{q}}$ must be singular on $\partial A$ unless $L\tilde{q}=0$ on $\partial A$. Moreover, to compute $L\tilde{q}$ requires computing $\Delta\tilde{q}$, which cannot be done efficiently by automatic differentiation when $\Omega$ is high-dimensional. Here, we derive a convenient alternative expression for the improper integral term

### v2

@@ 4.2 Selection: Estimating Relative Entropy Differences
[…]
Assume that $\tilde{m}$ and $\bar{m}$ are strictly positive. We define $\tilde{\mathds{P}}$ to be the law of the approximate transition path process corresponding to $\tilde{q}$ and $\tilde{m}$, i.e. the law of a weak solution of (8) with singular drift $2\varepsilon\nabla\log\tilde{q}(Y_{t})$ and initial condition $Y_{0}\sim\tilde{\eta}^{-1}m(x)\,\mathrm{d}S_{A}(x)$. We define $\bar{\mathds{P}}$ similarly, but with $\bar{q}$ and $\bar{m}$ in place of $\tilde{q}$ and $\tilde{m}$. We now explain how to estimate the relative entropy difference
@@
Suppose that we have a sample of $\tilde{N}$ paths $\tilde{Y}^{k}_{t}$ drawn from $\tilde{\mathds{P}}$:
@@
$\tilde{I}$ $:=\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\eta}\,\mathrm{d}S_{A}(x)$
$\qquad+\mathds{E}_{\mathds{P}_{\theta}}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$
[…]
@@
$\delta(\tilde{q},\tilde{m};\bar{q},\bar{m})=\log\left(\frac{\bar{\eta}}{\tilde{\eta}}\right)+\tilde{I}-\bar{I},$
so it remains to estimate the ratio $\frac{\bar{\eta}}{\tilde{\eta}}$. Analogous problems involving ratios of normalizing constants arise in Bayesian model selection and the calculation of alchemical free energy differences. Refer to [21] for a survey of methods for free energy calculations. We propose to use the Bennett Acceptance Ratio (BAR) method [4] to estimate $\frac{\bar{\eta}}{\tilde{\eta}}$. BAR is in some sense the optimal estimator of $\frac{\bar{\eta}}{\tilde{\eta}}$ given i.i.d. samples from $\tilde{m}\,\mathrm{d}S$ and $\bar{m}\,\mathrm{d}S$. More complex alternatives that involve sampling from multiple distributions, such as the Multistate Bennett Acceptance Ratio (MBAR) method, may be needed in cases where $\tilde{m}$ and $\bar{m}$ differ greatly [28].
Remark 4.4.
In practice, to sample from the approximate reactive flux densities $\tilde{m}$ and $\bar{m}$, one would most likely use a Markov chain Monte Carlo (MCMC) method. Many methods have been developed to sample distributions supported on submanifolds such as $\partial A$. For example, see [21, Sections 3.2.3-4]. Of course, MCMC would produce a correlated sample from the approximate reactive flux distribution, not an independent sample as assumed above.
4.3 Training On the Fly: Minimizing the Relative Entropy while Generating Reactive Trajectories
@@
(15) $m_{\theta}(x):=\eta_{\theta}^{-1}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
@@
$\eta_{\theta}:=\int_{\partial A}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x).$
We define $\mathds{P}_{\theta}$ to be the law of a weak solution of the stochastic differential equation (8) observed up to time $\tau$ for the approximate committor function $q_{\theta}$ and initial condition $Y_{0}\sim m_{\theta}$. In this section, we derive a consistent estimator of $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ given a sample from $\mathds{P}_{\theta}$. We then use the estimator in a stochastic gradient descent method to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ in Section 6.
For our results in this section, we require much stronger assumptions on $q_{\theta}$ and on the distribution of $\tau$ than those imposed elsewhere; cf. Assumption 4.3. The most significant of these new assumptions is that
(16) $Lq_{\theta}=0\text{ on }\partial A\text{ for all }\theta\in\mathbb{R}^{k}.$
Recall that for the exact committor $q$, or more precisely for the smooth extension guaranteed by Lemma 2.2, we have $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$; cf. Remark 2.4. In Section 5.1, we explain how to construct a practical families $q_{\theta}$ satisfying (16) when the boundary of $A$ is planar. However, we do not claim that (16) is necessary for differentiability. We simply cannot prove differentiability without imposing such an assumption.
@@
that appears in the change of measure formula (12) extends to a smooth function of $x$ defined on an open neighborhood of $\partial A$, even though we assume $q_{\theta}=0$ on $\partial A$. To see this, let $Lq_{\theta}$ take the place of $\tilde{q}$ in the proof of Lemma 3.1. Similarly,
@@
(17) $\mathds{E}_{\mathds{P}_{\theta}}[\exp(\gamma(\theta)\tau)]<\infty;$
that is, the moment generating function of $\tau$ under $\mathds{P}_{\theta}$ is finite on an open interval containing zero. We will not verify (17) for any particular family of approximate committor functions. However, we note that for the exact transition path process in one-dimension, under some rather stringent conditions on the potential energy $U$ and on the sets $A$ and $B$, one can show that $\tau$ converges in distribution as $\varepsilon\rightarrow 0$ to a shifted and scaled Gumbel random variable [7]. Under much more general conditions, escape times from basins of attraction of minima of $U$ are approximately exponentially distributed for the overdamped Langevin dynamics in the limit of small $\varepsilon$ [15]. The moment generating function of a Gumbel or exponential random variable will be finite on an open interval containing zero, but not globally finite, which motivates (17).
@@
    $M\geq\eta_{\theta}\geq\frac{1}{M}>0\text{ and }\lvert\nabla_{\theta}m_{\theta}\rvert\leq M$
@@
(18) $\mathds{E}_{\mathds{P}_{\theta}}[\exp(\gamma(\theta)\tau)]<\infty.$
Under Assumption 4.3, the relative entropy is differentiable.
Theorem 4.5.
Let Assumption 4.3 hold, and let $\mathds{P}_{\theta}$ and $m_{\theta}$ be defined as above. The relative entropy $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ is differentiable as a function of $\theta$, and
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ $=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\nabla_{\theta}\log q_{\theta}(Y_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y_{s})\,\mathrm{d}s,\right$.
@@
(19) $\begin{split}G_{N}&:=\frac{1}{N-1}\sum_{k=1}^{N}\left(\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-J_{N}\right)\\ &\qquad\qquad\qquad\times\left(\nabla_{\theta}\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-K_{N}\right).\end{split}$
Assumption 4.3 implies that $G_{N}$ has finite variance. We leave the proof to the reader, noting that it is similar to the proof of Lemma B.2. Using this gradient estimator, one can attempt to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ by stochastic gradient descent or variants such as Adam [19].
4.4 Importance Sampling and Direct Estimation of Relative Entropy
Using our change of measure formula, one could perhaps use importance sampling to estimate arbitrary expectations over the transition path process $\mathds{Q}$ given a sample of paths from the approximate process $\mathds{P}$. The discussion in this section is speculative. We do not claim that all estimators presented below are robust in practice or even that they would have finite variance under all conditions.
Let $g:\Omega\rightarrow\mathbb{R}$ be $\mathcal{F}_{\tau}$-measurable, i.e. a function of the transition path process observed up to time $\tau$. By (12), assuming for simplicity that $\tilde{q}=1$ on $\partial B$ and $m=\lvert\nabla\tilde{q}\rvert\exp(-U/\varepsilon)$, we have
$\mathds{E}_{\mathds{Q}}[g]$ $=\mathds{E}_{\mathds{P}}\left[Z_{\tau}^{-1}g\right]$
$=\frac{\eta}{\zeta}\mathds{E}_{\mathds{P}}\left[\exp\left(\int_{0}^{\tau(\omega)}\frac{L\tilde{q}}{\tilde{q}}(Y_{s}(\omega))\,\mathrm{d}s\right)g(\omega)\right],$
@@ 4.4 Importance Sampling and Direct Estimation of Relative Entropy
(20) $\bar{g}_{N}=\frac{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)g(Y^{k})}{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)},$
of $\mathds{E}_{\mathds{Q}}[g]$, where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. Experience with other estimators related to Girsanov’s theorem suggests that such a naïve estimator will not be robust. Some successes have been attained in the importance sampling of paths of diffusion processes by roughly similar means [34, 11, 30]. We hope that our results will facilitate the development of analogous methods for the transition path process. However, we note that [34, 11, 30] treat only paths observed up to a fixed finite time, not up to an unbounded stopping time as in our work. This is a significant challenge. In fact, since the stopping time is expected to have a moment generating function defined only a finite interval, it is not clear to us that the variance of $\bar{g}_{N}$ must be finite even if one assumes that $L\tilde{q}/\tilde{q}$ is bounded. By contrast, note that our estimators of relative entropy differences and gradients do have finite variance under plausible assumptions. Nonetheless, in Section 6, we present numerical experiments related to a simple problem for which the self-normalized importance sampling estimator produces good results.
@@
$\frac{\zeta}{\eta}$ $=\mathds{E}_{\mathds{P}}\left[\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)\right],$
@@
(21) $\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)$
for $\zeta/\eta$, where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. We do not expect this estimator to be any more robust than the more general importance sampling estimator above.
5 Numerical Methods
@@
5.1 Representing the Committor Function
@@
appearing in Assumption 4.3. In our approach, one must first choose a computable function $T:\mathscr{T}\rightarrow\mathbb{R}$ with the following properties:
  • $T$ extends to a smooth function on an open neighborhood of $\overline{\mathscr{T}}$.
@@
  • $T(x)>0$ for all $x\in\mathscr{T}\setminus\partial A$.
@@
where $T$ is as above and $w:\mathscr{T}\rightarrow\mathbb{R}$ belongs to some convenient family of smooth functions. Any $\tilde{q}$ of this form satisfies Assumption 3. In particular, the normal derivative of $\tilde{q}$ is positive over $\partial A$, since
$\nabla\tilde{q}(x)\cdot n(x)$ $=\big\{\nabla T(x)\exp(w(x))+\nabla w(x)T(x)\exp(w(x))\big\}\cdot n(x)$
@@
for $x\in\partial A$. Moreover, for any $T$ meeting the conditions above, the exact committor $q$ can be written in the form $q=T\exp(w)$ for $w=\log\left(q/T\right)$, and this $w$ extends to a smooth function on an open neighborhood of $\Omega$ by Lemma 3.1.
The condition $L\tilde{q}=0$ on $\partial A$ appearing in Assumption 4.3 is equivalent with a Neumann boundary condition on $w$. We have
(22) $L\tilde{q}=\left\{LT+2\varepsilon\nabla w\cdot\nabla T+T\left(Lw+\varepsilon\lvert\nabla w\rvert^{2}\right)\right\}\exp(w).$
@@
(23) $\nabla w(x)\cdot n(x)=-\frac{1}{2\varepsilon}\frac{LT(x)}{\lvert\nabla T(x)\rvert}$
for $x\in\partial A$. Of course, the difficulty of imposing such a boundary condition in practice depends on the geometry of $\partial A$.
5.2 Alternative Expression for the Improper Integral
Observe that $\frac{L\tilde{q}}{\tilde{q}}$ must be singular on $\partial A$ unless $L\tilde{q}=0$ on $\partial A$. Moreover, to compute $L\tilde{q}$ requires computing $\Delta\tilde{q}$. If $\tilde{q}$ is an artifical neural network and the state space is high-dimensional, then it may be inefficient to compute $\Delta\tilde{q}$ by automatic differentiation. See [16] for an explanation of the problem and for a general strategy that can be used to efficiently estimate $\Delta\tilde{q}$ in a stochastic gradient descent method at the cost of increasing the variance of the gradient estimates. Here, we derive a convenient alternative expression for the improper integral term